# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

### v2

@@ 2. Volume estimate
$\int_{B_{r}}\langle\nabla u,\nabla\phi\rangle\leq 0,\ \ \text{for any}\ \ 0\leq\phi\in C^{\infty}_{0}\left(B_{r}\right).$
@@
$\sup_{B_{\theta r}}u\leq C\left(\chi,s,\theta,\mathbb{S}_{\chi}(M)^{-1}\right)r^{-\frac{2\chi}{s(\chi-1)}}\left(\int_{B_{r}}(u^{+})^{s}\right)^{1/s}.$ (2.1)
@@
$\int_{B_{r}}\langle\nabla u,\nabla\phi\rangle\leq 0,\ \ \text{for any}\ \ 0\leq\phi\in W^{1,2}_{0}\left(B_{r}\right).$
For any $\eta(x)\in C_{0}^{\infty}\left(B_{r}\right)$, substituting $\phi=\eta^{2}u^{s-1}$ into the above inequality yields
$(s-1)\int_{B_{r}}|\nabla u|^{2}u^{s-2}\eta^{2}\leq-2\int_{B_{r}}\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle.$ (2.2)
@@
$\int_{B_{r}}|\nabla u|^{2}u^{s-2}\eta^{2}\leq\frac{4}{(s-1)^{2}}\int_{B_{r}}u^{s}|\nabla\eta|^{2}.$ (2.3)
@@
$\int_{B_{r}}|\nabla(\eta u^{s/2})|^{2}$ $=$ $\int\eta^{2}|\nabla u^{\frac{s}{2}}|^{2}+\int u^{s}|\nabla\eta|^{2}+s\int\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle$ (2.4)
@@
$\left(\int_{B_{r}}(\eta u^{s/2})^{2\chi}\right)^{\frac{1}{\chi}}\leq 10\mathbb{S}_{\chi}(M)^{-1}\int u^{s}|\nabla\eta|^{2}.$ (2.5)
@@
$|\nabla\eta_{k}|\leq\frac{2}{r_{k}-r_{k+1}}=\frac{2^{k+2}}{(1-\theta)r}.$ (2.6)
@@
$\|u\|_{L^{s_{k+1}}\left(B_{r_{k+1}}\right)}\leq\left\{\frac{160\times 4^{k}}{\mathbb{S}_{\chi}(M)(1-\theta)^{2}r^{2}}\right\}^{\frac{1}{s_{k}}}\|u\|_{L^{s_{k}}\left(B_{r_{k}}\right)}.$ (2.7)
@@
$\|u\|_{L^{s_{k+1}}\left(B_{r_{k+1}}\right)}\leq 4^{\sum_{i=0}^{k}\frac{i}{s_{i}}}\left\{\frac{160}{\mathbb{S}_{\chi}(M)(1-\theta)^{2}r^{2}}\right\}^{\sum_{i=0}^{k}\frac{1}{s_{i}}}\|u\|_{L^{s}\left(B_{r}\right)}.$ (2.8)
@@
$\|u\|_{L^{\infty}\left(B_{\theta r}\right)}\leq C\left(\chi,s,\mathbb{S}_{\chi}(M)^{-1}\right)[(1-\theta)r]^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$ (2.9)
@@
$\sup_{B_{\theta r}}u\leq\frac{2-s}{2}\sup_{B_{r}}u+\frac{s}{2}C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)^{\frac{2}{s}}[(1-\theta)r]^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$ (2.10)
Let $\tilde{s}=\theta r$, $t=r$, $\psi(s)=\sup_{B_{\theta r}}u$ and $\psi(t)=\sup_{B_{r}}u$. Then (2.10) can be rewritten as
$\psi(\tilde{s})\leq\frac{2-s}{2}\psi(t)+C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)(t-\tilde{s})^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
@@
Let $f(t)\geq 0$, $t\in[\tau_{0},\tau_{1}]$ with $\tau_{0}\geq 0$. Suppose for $\tau_{0}\leq t<\tilde{s}\leq\tau_{1}$,
$f(t)\leq\theta f(\tilde{s})+\frac{A}{(\tilde{s}-t)^{\alpha}}+B$
for some $\theta\in[0,1)$. Then for any $\tau_{0}\leq t<\tilde{s}\leq\tau_{1}$, there holds
$f(t)\leq c(\alpha,\theta)\left(\frac{A}{(\tilde{s}-t)^{\alpha}}+B\right).$
@@
$\frac{\left(\int_{M}|\nabla u|^{2}dv\right)^{\frac{1}{2}}}{\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}}$ $\geq$ $\sqrt{\mathbb{S}_{\chi}(M)}\frac{\left(\int_{M}u^{2\chi}dv\right)^{\frac{1}{2\chi}}}{\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}}$ (2.11)
@@
${\rm{vol}}\left(B_{x}(r)\right)\geq\left(r\sqrt{\mathbb{S}_{\chi}(M)}\right)^{2\alpha(m)}{2}^{-2\beta(m)}{\rm{vol}}\left(B_{x}\left(\frac{r}{2^{m}}\right)\right)^{\gamma(m)},$ (2.12)
@@
3.1. Linearization operator $\mathcal{L}$ of $p$-Laplacian.
@@
$\Delta_{p}u={\rm div}\left(|\nabla u|^{p-2}\nabla u\right).$ (3.1)
@@
$-\int_{\Omega}|\nabla v|^{p-2}\bm{\langle}\nabla v,\nabla\psi\bm{\rangle}+\int_{\Omega}av^{q}\psi=0.$
@@
On the other hand, it is easy to see from (3.1) that the linearization operator $\mathcal{L}$ of the $p$-Laplace operator is
$\mathcal{L}(\psi)={\rm div}\left(|\nabla u|^{p-2}A(\nabla\psi)\right),$
@@
$A(\nabla\psi)=\nabla\psi+(p-2)|\nabla u|^{-2}\bm{\langle}\nabla\psi,\nabla u\bm{\rangle}\nabla u.$
@@
Denote $f=|\nabla u|^{2}$. Then the linearization operator $\mathcal{L}$ of the $p$-Laplace operator can be rewritten as
$\mathcal{L}(\psi)={\rm div}\left(f^{p/2-1}A(\nabla\psi)\right),$ (3.3)
@@
$A(\nabla\psi)=\nabla\psi+(p-2)f^{-1}\bm{\langle}\nabla\psi,\nabla u\bm{\rangle}\nabla u.$ (3.4)
Next, we calculate the explicit expression $\mathcal{L}(f^{\alpha})$ for any $\alpha>0$ that will play a key role in our proof.
@@
$\begin{split}\mathcal{L}(f^{\alpha})=&\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+2\alpha f^{\alpha+\frac{p}{2}-2}\left(|\nabla\nabla u|^{2}+{\rm Ric}(\nabla u,\nabla u)\right)\\ &+\alpha(p-2)(\alpha-1)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}+2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle\end{split}$ (3.5)
@@
$\mathcal{L}(f^{\alpha})$ $=\alpha\text{div}\Big(f^{\alpha-1}f^{\frac{p}{2}-1}A(\nabla f)\Big)=\alpha\Big\langle\nabla(f^{\alpha-1}),f^{\frac{p}{2}-1}A(\nabla f)\Big\rangle+\alpha f^{\alpha-1}\mathcal{L}(f).$
@@
$\alpha\Big\langle\nabla(f^{\alpha-1}),f^{\frac{p}{2}-1}A(\nabla f)\Big\rangle=$ $\Big\langle\alpha(\alpha-1)f^{\alpha-2}\nabla f,f^{\frac{p}{2}-1}\nabla f+(p-2)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\nabla u\Big\rangle,$ (3.6)
@@
$\begin{split}\alpha f^{\alpha-1}\mathcal{L}(f)=&\alpha f^{\alpha-1}\Big(\left(\frac{p}{2}-1\right)f^{\frac{p}{2}-2}|\nabla f|^{2}+f^{\frac{p}{2}-1}\Delta f+(p-2)\left(\frac{p}{2}-2\right)f^{\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle^{2}\\ &+(p-2)f^{\frac{p}{2}-2}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle+(p-2)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\Delta u\Big).\end{split}$ (3.7)
@@
$\begin{split}\mathcal{L}(f^{\alpha})=&\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+\alpha f^{\alpha+\frac{p}{2}-2}\Delta f\\ &+\alpha(p-2)\left(\alpha+\frac{p}{2}-3\right)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}\\ &+\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle+\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle\Delta u.\end{split}$ (3.8)
@@
$\begin{split}\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle\Delta u&=2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle-2\alpha f^{\alpha+\frac{p}{2}-2}\langle\nabla\Delta u,\nabla u\rangle\\ &\quad-\alpha(p-2)\left(\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}\\ &\quad-\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle.\end{split}$ (3.9)
@@
$\langle\nabla\Delta u,\nabla u\rangle=\frac{1}{2}\Delta f-|\nabla\nabla u|^{2}-{\rm Ric}(\nabla u,\nabla u).$
@@
$\mathcal{L}(f^{\alpha})=$ $\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+2\alpha f^{\alpha+\frac{p}{2}-2}\left(|\nabla\nabla u|^{2}+{\rm Ric}(\nabla u,\nabla u)\right)$
@@
3.2. Precise estimate of $\mathcal{L}$.
In this subsection, we shall prove a precise estimate for $\mathcal{L}(f^{\alpha})$ when $v$ is a positive solution to equation (1.1).
@@
$\mathcal{L}(f^{\alpha})\geq 2\alpha f^{\alpha+\frac{p}{2}-2}\left(\beta_{n,p,q,\alpha}f^{2}-{\rm Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\right),$ (3.10)
@@
$u_{11}=\frac{1}{2}f^{-1}\bm{\langle}\nabla u,\nabla f\bm{\rangle}\ \ \ \text{and}\ \ \ \frac{|\nabla f|^{2}}{f}=4\sum_{i=1}^{n}u_{1i}^{2}.$ (3.11)
@@
$|\nabla\nabla u|^{2}\geq$ $\sum_{i=1}^{n}u_{1i}^{2}+\sum_{i=2}u_{ii}^{2}\geq\frac{|\nabla f|^{2}}{4f}+\frac{1}{n-1}\left(\sum_{i=2}u_{ii}\right)^{2}.$ (3.13)
@@
$\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}\left(f^{\alpha}\right)\geq&\frac{1}{2}\left(\alpha+\frac{p-3}{2}\right)\frac{|\nabla f|^{2}}{f}+\frac{1}{n-1}\left(\sum_{i=2}u_{ii}\right)^{2}+{\rm Ric}(\nabla u,\nabla u)\\ &+2(p-2)(\alpha-1)u_{11}^{2}+f^{1-\frac{p}{2}}\left(pf^{\frac{p}{2}}u_{11}+bce^{cu}f\right).\end{split}$ (3.14)
@@
$\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}\left(f^{\alpha}\right)\geq&2\left(\alpha+\frac{p-3}{2}\right)u_{11}^{2}+\frac{1}{n-1}\left(\sum_{i=2}^{n}u_{ii}\right)^{2}+{\rm Ric}(\nabla u,\nabla u)\\ &+2(p-2)(\alpha-1)u_{11}^{2}+f^{1-\frac{p}{2}}\left(pf^{\frac{p}{2}}u_{11}+bce^{cu}f\right).\end{split}$ (3.15)
@@
$\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}\left(f^{\alpha}\right)\geq&(p-1)(2\alpha-1)u_{11}^{2}-(n-1){\rm Ric}_{-}f+\left(p-\frac{2(p-1)}{n-1}\right)fu_{11}+\frac{f^{2}}{n-1}\\ &+b\left(c+\frac{2}{n-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{1}{n-1}\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}.\end{split}$ (3.16)
@@
$\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}(f^{\alpha})\geq&(p-1)(2\alpha-1)u_{11}^{2}-(n-1){\rm Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|+\frac{f^{2}}{n-1}\\ &+b\left(c+\frac{2}{n-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{1}{n-1}\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}.\end{split}$ (3.17)
@@
$\mathcal{L}(f^{\alpha})\geq$ $2\alpha f^{\alpha+\frac{p}{2}-2}\left(\frac{f^{2}}{n-1}-(n-1){\rm Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\right),$
@@
$\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}\mathcal{L}(f^{\alpha})\geq&(p-1)\left(2\alpha-1+\frac{p-1}{n-1}\right)u_{11}^{2}-(n-1){\rm Ric}_{-}f\\ &+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|+\frac{f^{2}}{n-1}\\ &+\frac{1}{n-1}\left(b^{2}e^{2cu}f^{2-p}-2(p-1)be^{cu}f^{1-\frac{p}{2}}u_{11}\right).\end{split}$ (3.18)
@@
$\begin{split}&(p-1)\left(2\alpha-1+\frac{p-1}{n-1}\right)u_{11}^{2}-2\frac{(p-1)}{n-1}be^{cu}f^{1-\frac{p}{2}}u_{11}\\ \geq&-\frac{(p-1)b^{2}e^{2cu}f^{2-p}}{((2\alpha-1)(n-1)+p-1)(n-1)}.\end{split}$ (3.19)

### v3

@@ 2. Volume estimate
$\int_{B_{r}}\langle\nabla u,\nabla\varphi\rangle\leq 0,\ \ \text{for any}\ \ 0\leq\varphi\in C^{\infty}_{0}\left(B_{r}\right).$
@@
(2.1) $\sup_{B_{\theta r}}u\leq C\left(\chi,s,\theta,\mathbb{S}_{\chi}(M)^{-1}\right)r^{-\frac{2\chi}{s(\chi-1)}}\left(\int_{B_{r}}(u^{+})^{s}\right)^{1/s}.$
@@
$\int_{B_{r}}\langle\nabla u,\nabla\varphi\rangle\leq 0,\ \ \text{for any}\ \ 0\leq\varphi\in W^{1,2}_{0}\left(B_{r}\right).$
For any $\eta(x)\in C_{0}^{\infty}\left(B_{r}\right)$, substituting $\varphi=\eta^{2}u^{s-1}$ into the above inequality yields
(2.2) $(s-1)\int_{B_{r}}|\nabla u|^{2}u^{s-2}\eta^{2}\leq-2\int_{B_{r}}\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle.$
@@
(2.3) $\int_{B_{r}}|\nabla u|^{2}u^{s-2}\eta^{2}\leq\frac{4}{(s-1)^{2}}\int_{B_{r}}u^{s}|\nabla\eta|^{2}.$
@@
(2.4) $\int_{B_{r}}|\nabla(\eta u^{s/2})|^{2}$ $=$ $\int\eta^{2}|\nabla u^{\frac{s}{2}}|^{2}+\int u^{s}|\nabla\eta|^{2}+s\int\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle$
@@
(2.5) $\left(\int_{B_{r}}(\eta u^{s/2})^{2\chi}\right)^{\frac{1}{\chi}}\leq 10\mathbb{S}_{\chi}(M)^{-1}\int u^{s}|\nabla\eta|^{2}.$
@@
(2.6) $|\nabla\eta_{k}|\leq\frac{2}{r_{k}-r_{k+1}}=\frac{2^{k+2}}{(1-\theta)r}.$
@@
(2.7) $\|u\|_{L^{s_{k+1}}\left(B_{r_{k+1}}\right)}\leq\left\{\frac{160\times 4^{k}}{\mathbb{S}_{\chi}(M)(1-\theta)^{2}r^{2}}\right\}^{\frac{1}{s_{k}}}\|u\|_{L^{s_{k}}\left(B_{r_{k}}\right)}.$
@@
(2.8) $\|u\|_{L^{s_{k+1}}\left(B_{r_{k+1}}\right)}\leq 4^{\sum_{i=0}^{k}\frac{i}{s_{i}}}\left\{\frac{160}{\mathbb{S}_{\chi}(M)(1-\theta)^{2}r^{2}}\right\}^{\sum_{i=0}^{k}\frac{1}{s_{i}}}\|u\|_{L^{s}\left(B_{r}\right)}.$
@@
(2.9) $\|u\|_{L^{\infty}\left(B_{\theta r}\right)}\leq C\left(\chi,s,\mathbb{S}_{\chi}(M)^{-1}\right)[(1-\theta)r]^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
@@
(2.10) $\sup_{B_{\theta r}}u\leq\frac{2-s}{2}\sup_{B_{r}}u+\frac{s}{2}C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)^{\frac{2}{s}}[(1-\theta)r]^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
Let $\widetilde{s}=\theta r$, $t=r$, $\psi(s)=\sup_{B_{\theta r}}u$ and $\psi(t)=\sup_{B_{r}}u$. Then (2.10) can be rewritten as
$\psi(\widetilde{s})\leq\frac{2-s}{2}\psi(t)+C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)(t-\widetilde{s})^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
@@
Let $f(t)\geq 0$, $t\in[\tau_{0},\tau_{1}]$ with $\tau_{0}\geq 0$. Suppose for $\tau_{0}\leq t<\widetilde{s}\leq\tau_{1}$,
$f(t)\leq\theta f(\widetilde{s})+\frac{A}{(\widetilde{s}-t)^{\alpha}}+B$
for some $\theta\in[0,1)$. Then for any $\tau_{0}\leq t<\widetilde{s}\leq\tau_{1}$, there holds
$f(t)\leq c(\alpha,\theta)\left(\frac{A}{(\widetilde{s}-t)^{\alpha}}+B\right).$
@@
(2.11) $\frac{\left(\int_{M}|\nabla u|^{2}dv\right)^{\frac{1}{2}}}{\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}}$ $\geq$ $\sqrt{\mathbb{S}_{\chi}(M)}\frac{\left(\int_{M}u^{2\chi}dv\right)^{\frac{1}{2\chi}}}{\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}}$
@@
(2.12) ${\rm{Vol}}\left(B_{x}(r)\right)\geq\left(r\sqrt{\mathbb{S}_{\chi}(M)}\right)^{2\alpha(m)}{2}^{-2\beta(m)}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2^{m}}\right)\right)^{\gamma(m)},$
@@
3.A. Linearization operator $\mathscr{L}$} of $p$-Laplacian.
@@
(3.1) $\Delta_{p}u={\rm div}\left(|\nabla u|^{p-2}\nabla u\right).$
@@
$-\int_{\Omega}|\nabla v|^{p-2}\langle\nabla v,\nabla\psi\rangle+\int_{\Omega}av^{q}\psi=0.$
@@
On the other hand, it is easy to see from (3.1) that the linearization operator $\mathscr{L}$} of the $p$-Laplace operator is
${\mathscr{L}}(\psi)={\rm div}\left(|\nabla u|^{p-2}A(\nabla\psi)\right),$
@@
$A(\nabla\psi)=\nabla\psi+(p-2)|\nabla u|^{-2}\langle\nabla\psi,\nabla u\rangle\nabla u.$
@@
Denote $f=|\nabla u|^{2}$. Then the linearization operator $\mathscr{L}$} of the $p$-Laplace operator can be rewritten as
(3.3) ${\mathscr{L}}(\psi)={\rm div}\left(f^{p/2-1}A(\nabla\psi)\right),$
@@
(3.4) $A(\nabla\psi)=\nabla\psi+(p-2)f^{-1}\langle\nabla\psi,\nabla u\rangle\nabla u.$
Next, we calculate the explicit expression ${\mathscr{L}}(f^{\alpha})$ for any $\alpha>0$ that will play a key role in our proof.
@@
(3.5) $\begin{split}{\mathscr{L}}(f^{\alpha})=&\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+2\alpha f^{\alpha+\frac{p}{2}-2}\left(|\nabla\nabla u|^{2}+\mathrm{Ric}(\nabla u,\nabla u)\right)\\ &+\alpha(p-2)(\alpha-1)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}+2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle\end{split}$
@@
${\mathscr{L}}(f^{\alpha})$ $=\alpha\text{div}\Big(f^{\alpha-1}f^{\frac{p}{2}-1}A(\nabla f)\Big)=\alpha\Big\langle\nabla(f^{\alpha-1}),f^{\frac{p}{2}-1}A(\nabla f)\Big\rangle+\alpha f^{\alpha-1}\mathcal{L}(f).$
@@
(3.6) $\alpha\Big\langle\nabla(f^{\alpha-1}),f^{\frac{p}{2}-1}A(\nabla f)\Big\rangle=$ $\Big\langle\alpha(\alpha-1)f^{\alpha-2}\nabla f,f^{\frac{p}{2}-1}\nabla f+(p-2)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\nabla u\Big\rangle,$
@@
(3.7) $\begin{split}\alpha f^{\alpha-1}{\mathscr{L}}(f)=&\alpha f^{\alpha-1}\Big(\left(\frac{p}{2}-1\right)f^{\frac{p}{2}-2}|\nabla f|^{2}+f^{\frac{p}{2}-1}\Delta f+(p-2)\left(\frac{p}{2}-2\right)f^{\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle^{2}\\ &+(p-2)f^{\frac{p}{2}-2}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle+(p-2)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\Delta u\Big).\end{split}$
@@
(3.8) $\begin{split}{\mathscr{L}}(f^{\alpha})=&\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+\alpha f^{\alpha+\frac{p}{2}-2}\Delta f\\ &+\alpha(p-2)\left(\alpha+\frac{p}{2}-3\right)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}\\ &+\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle+\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle\Delta u.\end{split}$
@@
(3.9) $\begin{split}\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle\Delta u&=2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle-2\alpha f^{\alpha+\frac{p}{2}-2}\langle\nabla\Delta u,\nabla u\rangle\\ &\quad-\alpha(p-2)\left(\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}\\ &\quad-\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle.\end{split}$
@@
$\langle\nabla\Delta u,\nabla u\rangle=\frac{1}{2}\Delta f-|\nabla\nabla u|^{2}-\mathrm{Ric}(\nabla u,\nabla u).$
@@
${\mathscr{L}}(f^{\alpha})=$ $\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+2\alpha f^{\alpha+\frac{p}{2}-2}\left(|\nabla\nabla u|^{2}+\mathrm{Ric}(\nabla u,\nabla u)\right)$
@@
3.B. Precise estimate of $\mathscr{L}$}.
In this subsection, we shall prove a precise estimate for ${\mathscr{L}}(f^{\alpha})$ when $v$ is a positive solution to equation (1.1).
@@
(3.10) ${\mathscr{L}}(f^{\alpha})\geq 2\alpha f^{\alpha+\frac{p}{2}-2}\left(\beta_{n,p,q,\alpha}f^{2}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\right),$
@@
(3.11) $u_{11}=\frac{1}{2}f^{-1}\langle\nabla u,\nabla f\rangle\ \ \ \text{and}\ \ \ \frac{|\nabla f|^{2}}{f}=4\sum_{i=1}^{n}u_{1i}^{2}.$
@@
(3.13) $|\nabla\nabla u|^{2}\geq$ $\sum_{i=1}^{n}u_{1i}^{2}+\sum_{i=2}u_{ii}^{2}\geq\frac{|\nabla f|^{2}}{4f}+\frac{1}{n-1}\left(\sum_{i=2}u_{ii}\right)^{2}.$
@@
(3.14) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}\left(f^{\alpha}\right)\geq&\frac{1}{2}\left(\alpha+\frac{p-3}{2}\right)\frac{|\nabla f|^{2}}{f}+\frac{1}{n-1}\left(\sum_{i=2}u_{ii}\right)^{2}+\mathrm{Ric}(\nabla u,\nabla u)\\ &+2(p-2)(\alpha-1)u_{11}^{2}+f^{1-\frac{p}{2}}\left(pf^{\frac{p}{2}}u_{11}+bce^{cu}f\right).\end{split}$
@@
(3.15) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}\left(f^{\alpha}\right)\geq&2\left(\alpha+\frac{p-3}{2}\right)u_{11}^{2}+\frac{1}{n-1}\left(\sum_{i=2}^{n}u_{ii}\right)^{2}+\mathrm{Ric}(\nabla u,\nabla u)\\ &+2(p-2)(\alpha-1)u_{11}^{2}+f^{1-\frac{p}{2}}\left(pf^{\frac{p}{2}}u_{11}+bce^{cu}f\right).\end{split}$
@@
(3.16) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}\left(f^{\alpha}\right)\geq&(p-1)(2\alpha-1)u_{11}^{2}-\mathrm{Ric}_{-}f+\left(p-\frac{2(p-1)}{n-1}\right)fu_{11}+\frac{f^{2}}{n-1}\\ &+b\left(c+\frac{2}{n-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{1}{n-1}\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}.\end{split}$
@@
(3.17) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq&(p-1)(2\alpha-1)u_{11}^{2}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|+\frac{f^{2}}{n-1}\\ &+b\left(c+\frac{2}{n-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{1}{n-1}\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}.\end{split}$
@@
${\mathscr{L}}(f^{\alpha})\geq$ $2\alpha f^{\alpha+\frac{p}{2}-2}\left(\frac{f^{2}}{n-1}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\right),$
@@
(3.18) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq&(p-1)\left(2\alpha-1+\frac{p-1}{n-1}\right)u_{11}^{2}-\mathrm{Ric}_{-}f\\ &+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|+\frac{f^{2}}{n-1}\\ &+\frac{1}{n-1}\left(b^{2}e^{2cu}f^{2-p}-2(p-1)be^{cu}f^{1-\frac{p}{2}}u_{11}\right).\end{split}$
@@
(3.19) $\begin{split}&(p-1)\left(2\alpha-1+\frac{p-1}{n-1}\right)u_{11}^{2}-2\frac{(p-1)}{n-1}be^{cu}f^{1-\frac{p}{2}}u_{11}\\ \geq&-\frac{(p-1)b^{2}e^{2cu}f^{2-p}}{((2\alpha-1)(n-1)+p-1)(n-1)}.\end{split}$